Cory's age is 20% of her father's age. Forty-two years from now, Cory's age will be two-thirds (2/3) of her father's age. How old are Cory and her father now?
Answer by mananth(16946)

cory's age = 20% of father's age
so father is 5 times cory's age
let cory's age = x
father's age = 5x
42 years later
Cory's age x+42
father's age 5x+42
If Cory 's age = 2/3 of father's age then father is 1.5 times cory's age
father's age 1.5 times older
5x+42=1.5(x+42)
5x+42=1.5x+63
3.5x=63 -42
3.5x=21
/3.5
x=6 Cory's age
x+20=26 years father's age